OpenStudy (anonymous):

in country a about 5 times as many cars are manufactured per day than in country b. if the total number of these cars manufactured per day is 36,636, find the number manufactured in country b and in a

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

According to the question a = 5b ----(1 and a+b = 36,636 ----(2 from (1 and (2 we have. 5b + b = 36,636 6b = 36,636 \[b= \frac{ 36636}{6} = 6106\] a = 5b = 5*6106=30530 Hence country a manufactures 30530 cars and country b manufacture 6106 cars. @crystalG

OpenStudy (anonymous):

omg thank u soooo much!

OpenStudy (anonymous):

@crystalG Medal please...